OpenStudy (anonymous):

can someone give me a refresh on how to add,subtract,multiply and divide fractions,

OpenStudy (anonymous):

Sure, adding first, to add fractions they have to have the same denominatore Ex. 1/4 + 2/5 you cant add at its state but if you make it 5/20 + 8/20 you can add it

OpenStudy (anonymous):

ahh ok gotcha and then what if the answer to this problem is a whole number? just reviewing what I can remember from a college test called the pert I have to retake

OpenStudy (anonymous):

i feel so dumb with the little things i forgot

OpenStudy (anonymous):

you mean like 20/20 or do you mean like 23/20?

OpenStudy (anonymous):

like in the question i remember it had me add to fractions but the answers where whole numbers

OpenStudy (anonymous):

so like 1 and 1/3?

OpenStudy (ranga):

Two fractions can add to become a whole number: 1/5 + 4/5 = 5/5 = 1

OpenStudy (anonymous):

i think i get it mainly i just forgot you had to find a common denominator when adding ok and subtracting?

OpenStudy (anonymous):

ok anything else?

OpenStudy (anonymous):

yes what do i do when subtracting is it the same thing?

OpenStudy (ranga):

subtraction: same rule as addition. find common denominator.

OpenStudy (anonymous):

ok and multiply you switch and add correct?

OpenStudy (anonymous):

Yup :P thx ranga

OpenStudy (ranga):

multiplication: multiply numerator and leave it on top. multiply denominator leave it at bottom. simplify.

OpenStudy (the_fizicx99):

Multiplication and division don't require the same denominator

OpenStudy (anonymous):

so i just multiply straight across?

OpenStudy (ranga):

yes. division. flip the bottom and multiply.

OpenStudy (anonymous):

ok flip the bottom on both ? and multiply straight across?

OpenStudy (anonymous):

for division

OpenStudy (the_fizicx99):

For division; keep the first fraction, switch the division into a multiplication sign and switch the second fraction. (recipricol)

OpenStudy (the_fizicx99):

Flip*

OpenStudy (ranga):

no. a/b divide by c/d = a/b * d/c = ad/bc

OpenStudy (the_fizicx99):

Yes like @ranga said :D

OpenStudy (anonymous):

ahh ok so i keep first one flip second one and multiply for dividing fracrtions?

OpenStudy (the_fizicx99):

Yes

OpenStudy (anonymous):

and multiply i just multiply straight across?

OpenStudy (the_fizicx99):

Yup

OpenStudy (ranga):

yes. 2/3 * 5/6 = (2*5) / (3*6) = 10/18 = 5/9

OpenStudy (anonymous):

ok sweet i wish i could give both medal but i do have a couple more questions so look out for them if you know thank you so much guys:)

OpenStudy (ranga):

yw.

OpenStudy (the_fizicx99):

Np. Gl

OpenStudy (anonymous):

you just ask it here if you l try to help

OpenStudy (anonymous):

*you can just typt here if you want ill try to help